A cheetah can run at a maximum speed
92 km/h and a gazelle can run at a maximum speed of 72.7 km/h.
If both animals are running at full speed,
with the gazelle 72.9 m ahead, how long before
the cheetah catches its prey?
Answer in units of s.
1
Expert's answer
2019-10-09T10:41:09-0400

Speeds of animals

"v_c=92\:\rm km/h=25.6 \:\rm m/s"

"v_g=72.7\:\rm km/h=20.2 \:\rm m/s"

The relative speed of the cheetah with respect to gazelle


"v=v_c-v_g=25.6\:\rm m/s-20.2\: m/s=5.4\: m/s"


Therefore, the time when the cheetah catches its prey


"t=\frac{d}{v}=\frac{72.9\:\rm m}{5.4\:\rm m/s}=13.5\:\rm s"